# Randomized Space–Time Coded Stacked Intelligent Metasurfaces for Massive Multiuser Downlink Connectivity

### v1

@@ II-A Wave propagation model through the ST-SIM
We assume that the S-coded block of the ST-SIM, including the layers $\ell\in\mathcal{L}^{\text{(s)}}\triangleq\{2,3,\ldots,L\},$ is implemented using amplitude-controlled (AC) and/or phase-controlled (PC) metasurfaces [18]. AC layers allow programmable amplitude control via active circuitry, whereas PC layers, which are nearly passive, permit programmable phase shifts. We denote the corresponding index sets with $\mathcal{L}^{\text{(s)}}_{\text{ac}}$ and $\mathcal{L}^{\text{(s)}}_{\text{pc}}$ such that
@@
[…]
Regarding the ST block (i.e., layer $1$), all meta-atoms are phase-controlled. Let $\delta_{z}(t)$ denote the time-varying transmission coefficient of the $z$-th meta-atom in the initial ST-coded DAL, for $t\in[0,T)$, where the one-dimensional index $z$ is obtained from the 2-D coordinates $(z_{x},z_{y})$ as
$\delta_{z}(t)=\sum_{m=0}^{M-1}\delta^{(m)}_{z}\,p(t-m\,T_{\text{s}})$ (5)
where $p(t)=\Pi\!\left(\frac{t-T_{\text{s}}/2}{T_{\text{s}}}\right)$ represents a rectangular pulse of duration $T_{\text{s}}$, and $\delta^{(m)}_{z}=\beta\,e^{j\psi^{(m)}_{z}}$ is the transmission coefficient of the $z$-th PC meta-atom in the $m$-th time slot, for $m\in\mathcal{M}\triangleq\{0,1,\ldots,M-1\}$, with digitally controllable phase $\psi^{(m)}_{z}$ and fixed amplitude $\beta\leq 1$. Let ${\mathbf{\Delta}}(t)\triangleq\mathrm{diag}\,[\boldsymbol{\delta}(t)]$ denote the diagonal matrix whose main diagonal collects the rapidly TV coefficients $\boldsymbol{\delta}(t)\triangleq[\delta_{0}(t),\ldots,\delta_{Z-1}(t)]^{\top}\in\mathbb{C}^{Z}$ of the ST-coded initial layer.
TABLE I: Main system parameters.
@@
$d_{\text{meta}}$ Spacing between adjacent meta-atoms $Z$ Number of meta-atoms of the input ST-coded DAL
$s_{\text{lay}}$ Spacing between adjacent layers of the ST-SIM $Q$ Number of meta-atoms in the intermediate S-coded layers
@@
$T_{s}$ Duration of each time-slot ($T/M$) $V$ Number of meta-atoms in the terminal S-coded DAL
Regarding the wave propagation between successive layers, all forward-propagation matrices share the same elementwise expression derived from the Rayleigh–Sommerfeld diffraction equation [12, 3]:
@@
where $\kappa_{0}\triangleq 2\pi/\lambda_{0}$ is the wave number, with wavelength $\lambda_{0}=c/f_{0}$ and light speed (in the vacuum) $c=3\cdot 10^{8}$ m/s, $A$ is the effective area of the transmitting element, $s$ is the inter-layer separation, and $d$ is the propagation distance between the two elements. Specifically, propagation between the UPA and the ST-coded DAL (layer $1$) is described by the following matrix
@@
Similarly, propagation between the ST-coded DAL (layer $1$) and layer $2$ is described by
@@
Finally, propagation between layer L1L-1 and the terminal S-coded DAL (layer $L$) is described by
@@
represents the distance between the $q$-th meta-atom of layer L1L-1 and the $v$-th transmitting meta-atom of the terminal DAL. The end-to-end forward propagation TV matrix across the ST-SIM, from the input of the ST-coded DAL to the output of the S-coded DAL, is given by
@@
${\mathbf{G}}_{0}\triangleq{\mathbf{\Gamma}}_{L}\,{\mathbf{W}}_{L}\,{\mathbf{\Gamma}}_{L-1}\,{\mathbf{W}}_{L-1}\cdots{\mathbf{\Gamma}}_{2}\,{\mathbf{W}}_{2}\in\mathbb{C}^{V\times Z}$ (16)
models the forward propagation through the S-coded multi-layer block, and ${\mathbf{\Delta}}(t)$ is the TV diagonal response of the ST-coded initial DAL. For clarity, the main system parameters are summarized in Table I, whereas descriptions and dimensions of all matrices characterizing the ST-SIM are listed in Table II.
TABLE II: Dimensions of the main matrices in the ST–SIM model.
Symbol Description Dimensions
${\mathbf{W}}_{1}$ UPA $\to$ ST-coded DAL (layer 1) $Z\times N$
${\mathbf{W}}_{2}$ ST-coded DAL (layer 1) $\to$ layer 2 $Q\times Z$
${\mathbf{W}}_{L}$ layer $L\!-\!1\to$ S-coded DAL (layer $L$) $V\times Q$
[…]
${\mathbf{\Gamma}}_{\text{dal}}$ S-coded DAL coefficients $V\times V$
${\mathbf{\Delta}}(t)$ ST-coded DAL response (layer 1) $Z\times Z$
${\mathbf{G}}_{0}$ S-coded block response (ST-coded DAL output $\to$ S-coded DAL output) $V\times Z$
${\mathbf{G}}(t)$ ST–SIM response (ST-coded DAL input $\to$ S-coded DAL output) $V\times Z$
$\widetilde{{\mathbf{G}}}(t)\triangleq{\mathbf{G}}(t){\mathbf{W}}_{1}$ UPA output $\to$ S-coded DAL output $V\times N$
@@ III Signal models
$=\beta^{2}\,\frac{\mathscr{E}}{T_{\mathrm{b}}}\,\sum_{n=0}^{N-1}\sum_{z=0}^{Z-1}\big|[{\mathbf{W}}_{1}]_{z,n}\big|^{2}\,\|{\mathbf{g}}_{z}^{(0)}\|^{2}$ (24)
where $\langle\cdot\rangle$ denotes the time-average operator, $\|\cdot\|$ is the Frobenius matrix norm, and ${\mathbf{G}}_{0}=[{\mathbf{g}}_{0}^{(0)},\ldots,{\mathbf{g}}_{Z-1}^{(0)}]$. The last two equalities follow from the statistical independence among the data streams $\{d_{n}(t)\}_{n=0}^{N-1}$ and the SIM transmission coefficients $\{\delta_{n}(t)\}_{n=0}^{N-1}$, together with the pairwise uncorrelated nature of their components. Hereinafter, we will assume that no power amplification is realized by the SIM and, then, we will enforce the following constraint in its design:
$\|{\mathbf{g}}_{z}^{(0)}\|^{2}=\frac{1}{\beta^{2}\,\|{\mathbf{W}}_{1}\|^{2}}\>,\quad\text{for $z\in\mathcal{Z}$}$ (25)
which entails that the radiated power is equal to $\mathscr{P}_{\text{rad}}={\mathscr{E}}/{T_{\text{b}}}$. Strictly speaking, the AC layers in the S-coded part of the SIM are used only to compensate for the propagation losses inside the SIM. This choice differs from [12], where the complex transmission coefficients of each layer exhibit unit amplitude implying, thus, that each layer can be realized without utilizing resistive components (local design). We, instead, enlarge the set of feasible solutions of our design problem by allowing that the globally radiated power is preserved [18].
@@ IV User scheduling with partial CSIT
and feeds the information $\text{SINR}_{u,n^{\star}_{u,m}}(m)$ back to the BS, for each time-slot. Based on the partial information provided by all the $U$ users in the network, the base station opportunistically schedules transmissions in every time-slot to the $N$ users who exhibit the highest $\text{SINR}_{u,n^{\star}_{u,m}}(m)$ values, where the maximization is performed over $u$ for each steering vector index $n\in\mathcal{N}$. In this manner, the system can opportunistically serve up to NMN\,M different users in each time interval $[0,T)$. It is worth noting that this partial CSI can be fed back to the BS using either an analog strategy, i.e., transmitting an unquantized scalar, or a digital strategy, that is transmitting a quantized version of $\text{SINR}_{u,n^{\star}_{u,m}}(m)$.
We now assume that, in each time-slot $m\in\mathcal{M}$ and for every steering vector index $n\in\mathcal{N}$, there exists at least one user ${u}^{\star}_{n}\in\mathcal{U}$ for which the SINR is maximum when $n$ is fixed. We define the set ${\mathcal{U}}^{\star}\triangleq\{{u}^{\star}_{0},{u}^{\star}_{1},\ldots,{u}^{\star}_{N-1}\}\subset\mathcal{U}$ that collects these $N$ selected users, and denote the corresponding maximum SINR values concisely as $\text{SINR}^{\star}_{{u}^{\star}_{n}}(m)$. Given the received signal model (29), the achievable rate for user ${u}^{\star}_{n}$ in the $m$-th time-slot is
@@
where $\mathcal{R}_{{u}^{\star}_{n}}(m)$ denotes the achievable rate for user ${u}^{\star}_{n}$ in the $m$-th time-slot and is given by (34).
@@
Remark 1: Unlike conventional S-coded SIM schemes, the time modulation employed in the proposed ST-SIM architecture enables the exploitation of multiuser diversity. When the channel between the SIM and the users is slowly time-varying, with a coherence time $T$ spanning multiple symbol periods (i.e., $T=M\,T_{s}=M\,P\,T_{\text{b}}$, with $M\,P\gg 1$), the channel remains nearly constant over several transmission intervals. In such a scenario, only $N$ out of the $U$ users can be scheduled and effectively served within each coherence interval. As the total number of users $U$ increases, this leads to significant unfairness across the network [29]. In contrast, the proposed ST-SIM-based beamforming approach introduces artificial time variations into the wireless channel via time modulation. This temporal diversity allows the system to opportunistically schedule up to $MN$ users for transmission. Consequently, the ST-SIM architecture improves fairness and enables the exploitation of multiuser diversity, even under slow channel dynamics.
@@ V Transmission Coefficient Synthesis for randomized ST-SIM
At the beginning of each time slot, the ST-SIM has to generate a set of $N$ beamforming vectors $\{\widetilde{\boldsymbol{g}}_{n}(t)\}_{n=0}^{N-1}$, each used to modulate the corresponding information signal $d_{n}(t)$. Recalling the structure of the overall forward propagation matrix in (15), the generation of these steering vectors involves two distinct steps. The first step consists of generating a fixed matrix ${\mathbf{G}}_{0}$, which models the forward propagation through the S-coded multilayer block of the SIM. This step is performed once, at the beginning of each channel coherence interval of duration $T$. The second step, in contrast, is repeated at the beginning of each time-slot of duration $T_{\text{s}}$ and consists of randomly generating the rapidly TV transmission coefficients $\boldsymbol{\delta}(t)$ of the initial ST-coded DAL.
Step 1 - Generation of the matrix ${\mathbf{G}}_{0}$: Let ${\mathbf{G}}_{\text{targ}}\in\mathbb{C}^{V\times Z}$ denote the target matrix that models the forward propagation through the S-coded multilayer block of the SIM, namely, from the output of the ST-coded DAL to the output of the S-coded DAL. It can be expressed as
@@
where ${\mathbf{g}}_{\text{targ},n}\in\mathbb{C}^{V}$ denotes the desired steering vector associated with the $n$-th transmitted stream. The columns of ${\mathbf{G}}_{\text{targ}}$ are orthogonal, i.e., ${\mathbf{g}}_{\text{targ},z_{1}}]^{\mathsf{H}}\,{\mathbf{g}}_{\text{targ},z_{2}}=0$ for $z_{1}\neq z_{2}\in\mathcal{Z}$, and obey the norm constraint (25).
Our objective is to synthesize the transmission coefficients $\{\boldsymbol{\gamma}_{\ell}\}_{\ell\in\mathcal{L}^{(\text{s})}}$ of the S-coded metasurface layers such that the matrix ${\mathbf{G}}_{0}$ in (15) closely approximates ${\mathbf{G}}_{\text{targ}}$. This leads to the constrained least-squares (LS) optimization problem
@@
$\|{\mathbf{g}}_{z}^{(0)}\|^{2}=\frac{1}{\beta^{2}\,\|{\mathbf{W}}_{1}\|^{2}},\quad z\in\mathcal{Z}$ (44)
@@
The PGD iterations (49)-(50) continue until a convergence criterion is met, either a sufficiently small variation in the cost function or a maximum number of iterations $\kappa_{\max}$ is reached. Upon convergence, the set of optimized transmission coefficients $\{\boldsymbol{\gamma}^{\star}_{\ell}\}_{\ell\in\mathcal{L}^{(\text{s})}}$ configures the S-coded part of the SIM.
@@
with ${\mathbf{1}}_{Q}$ denoting the $Q\times 1$ all-ones vector, and $\circ$ being the Hadamard (element-wise) product.

### v2

@@ II-A Wave propagation model through the ST-SIM
We assume that the S-only block of the ST-SIM, including the layers $\ell\in\mathcal{L}^{\text{(s)}}\triangleq\{2,3,\ldots,L\},$ is implemented using amplitude-controlled (AC) and/or phase-controlled (PC) metasurfaces [18]. AC layers allow programmable amplitude control via active circuitry, whereas PC layers, which are nearly passive, permit programmable phase shifts. We denote the corresponding index sets with $\mathcal{L}^{\text{(s)}}_{\text{ac}}$ and $\mathcal{L}^{\text{(s)}}_{\text{pc}}$ such that
@@
Let $\gamma_{\ell,q}=\alpha_{\ell,q}\,e^{j\phi_{\ell,q}}$, for $q\in\mathcal{Q}$ and $\ell\in\mathcal{L}^{\text{(s)}}-\{L\}$, denote the transmission coefficient of the $q$-th meta-atom in the $\ell$-th layer of the S-only block (excluding the terminal DAL), where $\alpha_{\ell,q}$ and $\phi_{\ell,q}$ are its amplitude and phase, respectively. We define $\boldsymbol{\gamma}_{\ell}\triangleq[\gamma_{\ell,0},\ldots,\gamma_{\ell,Q-1}]^{\top}\in\mathbb{C}^{Q}$, which collects the transmission coefficients of all $Q$ meta-atoms in the $\ell$-th layer, and the diagonal matrix ${\mathbf{\Gamma}}_{\ell}\triangleq\mathrm{diag}(\boldsymbol{\gamma}_{\ell})$.
TABLE II: Main system parameters.
@@
$d_{\text{meta}}$ Spacing between adjacent meta-atoms $Z$ Number of meta-atoms of the input ST DAL
$s_{\text{lay}}$ Spacing between adjacent layers of the ST-SIM $Q$ Number of meta-atoms in the intermediate S-only layers
@@
$T_{s}$ Duration of each time-slot ($T/M$) $V$ Number of meta-atoms in the terminal S-only DAL
[…]
Regarding the ST block (i.e., layer $1$), all meta-atoms are phase-controlled. Let $\delta_{z}(t)$ denote the time-varying transmission coefficient of the $z$-th meta-atom in the initial ST DAL, for $t\in[0,T)$, where the one-dimensional index $z$ is obtained from the 2-D coordinates $(z_{x},z_{y})$ as
$\delta_{z}(t)=\sum_{m=0}^{M-1}\delta^{(m)}_{z}\,p(t-m\,T_{\text{s}})$ (5)
where $p(t)=\Pi\!\left(\frac{t-T_{\text{s}}/2}{T_{\text{s}}}\right)$ represents a rectangular pulse of duration $T_{\text{s}}$, and $\delta^{(m)}_{z}=\beta\,e^{j\psi^{(m)}_{z}}$ is the transmission coefficient of the $z$-th PC meta-atom in the $m$-th time slot, for $m\in\mathcal{M}\triangleq\{0,1,\ldots,M-1\}$, with digitally controllable phase $\psi^{(m)}_{z}$ and fixed amplitude $\beta\leq 1$. Let ${\mathbf{\Delta}}(t)\triangleq\mathrm{diag}\,[\boldsymbol{\delta}(t)]$ denote the diagonal matrix whose main diagonal collects the rapidly TV coefficients $\boldsymbol{\delta}(t)\triangleq[\delta_{0}(t),\ldots,\delta_{Z-1}(t)]^{\top}\in\mathbb{C}^{Z}$ of the ST initial layer.
Regarding the wave propagation between successive layers, all forward-propagation matrices share the same elementwise expression derived from the Rayleigh–Sommerfeld diffraction propagator [12, 3]:
@@
where $\kappa_{0}\triangleq 2\pi/\lambda_{0}$ is the free-space wave number, with corresponding wavelength $\lambda_{0}=c/f_{0}$ and light speed $c=3\cdot 10^{8}$ m/s.
TABLE III: Dimensions of the main matrices in the ST–SIM model.
Symbol Description Dimensions
${\mathbf{W}}_{1}$ UPA $\to$ ST-DAL (layer 1) $Z\times N$
${\mathbf{W}}_{2}$ ST DAL (layer 1) $\to$ layer 2 $Q\times Z$
${\mathbf{W}}_{L}$ layer $L\!-\!1\to$ S-only DAL (layer $L$) $V\times Q$
[…]
${\mathbf{\Gamma}}_{\text{dal}}$ S-only DAL coefficients $V\times V$
${\mathbf{\Delta}}(t)$ ST DAL response (layer 1) $Z\times Z$
${\mathbf{G}}_{0}$ S-only block response (ST DAL output $\to$ S-only DAL output) $V\times Z$
${\mathbf{G}}(t)$ ST-SIM response (ST-DAL input $\to$ S-only DAL output) $V\times Z$
$\widetilde{{\mathbf{G}}}(t)\triangleq{\mathbf{G}}(t){\mathbf{W}}_{1}$ UPA output $\to$ S-only DAL output $V\times N$
It models free-space EM wave propagation between two radiating elements separated by distance $d$, with effective aperture area $A$, and inter-layer separation $s$. For a given metasurface, the transmission coefficients are typically obtained from unit-cell full-wave EM simulations under the assumption of local periodicity. This approximation is commonly adopted in metasurface modeling and is valid when higher-order grating modes are evanescent and sufficiently attenuated before reaching adjacent metasurface layers. Under these conditions, the interaction between layers can be accurately described through effective transmission coefficients and diffraction-based propagation models. Although the detailed EM implementation of the unit cells is beyond the scope of this work, the inter-element spacing and inter-layer separations considered in our examples in Section VI are consistent with regimes in which the local periodicity assumption provides a good approximation of the inter-layer propagation. Specifically, propagation between the UPA and the ST DAL (layer $1$) is described by the following matrix
@@
Similarly, propagation between the ST DAL (layer $1$) and layer $2$ is described by
@@
Finally, propagation between layer L1L-1 and the terminal S-only DAL (layer $L$) is described by
@@
represents the distance between the $q$-th meta-atom of layer L1L-1 and the $v$-th transmitting meta-atom of the terminal DAL. The end-to-end forward propagation TV matrix across the ST-SIM, from the input of the ST DAL to the output of the S-only DAL, is given by
@@
${\mathbf{G}}_{0}\triangleq{\mathbf{\Gamma}}_{\mathrm{dal}}\,{\mathbf{W}}_{L}\,{\mathbf{\Gamma}}_{L-1}\,{\mathbf{W}}_{L-1}\cdots{\mathbf{\Gamma}}_{2}\,{\mathbf{W}}_{2}\in\mathbb{C}^{V\times Z}$ (16)
models the forward propagation through the S-only multi-layer block, and ${\mathbf{\Delta}}(t)$ is the TV diagonal response of the ST initial DAL. For clarity, the main system parameters are summarized in Table II, whereas descriptions and dimensions of all matrices characterizing the ST-SIM are listed in Table III.
@@ III Signal models
$=\mathscr{P}_{\text{sig}}\,\frac{\beta^{2}}{N}\,\sum_{n=0}^{N-1}\sum_{z=0}^{Z-1}\big|[{\mathbf{W}}_{1}]_{z,n}\big|^{2}\,\|{\mathbf{g}}_{z}^{(0)}\|^{2}$ (24)
where $\langle\cdot\rangle$ denotes the time-average operator, $\|\cdot\|$ is the Frobenius matrix norm, ${\mathbf{G}}_{0}=[{\mathbf{g}}_{0}^{(0)},\ldots,{\mathbf{g}}_{Z-1}^{(0)}]$, and $\mathscr{P}_{\text{sig}}\triangleq\tfrac{N\mathscr{E}}{T_{\mathrm{b}}}$ denotes the average power of the input signal ${\mathbf{d}}(t)$. The last two equalities follow from the assumption that the data streams $\{d_{n}(t)\}_{n=0}^{N-1}$ are i.i.d. and the SIM transmission coefficients $\{\delta_{n}(t)\}_{n=0}^{N-1}$, together with the pairwise uncorrelated nature of their components. Hereinafter, we will assume that no power amplification is realized by the SIM and, then, we will enforce the following constraint in its design:
$\|{\mathbf{g}}_{z}^{(0)}\|^{2}=\frac{N}{\beta^{2}\,\|{\mathbf{W}}_{1}\|^{2}}\>,\quad\text{for $z\in\mathcal{Z}$}$ (25)
which entails that the radiated power is equal to $\mathscr{P}_{\text{rad}}=\mathscr{P}_{\text{sig}}$. Strictly speaking, the AC layers in the S-only part of the SIM are used only to compensate for the propagation losses inside the SIM. This choice differs from [12], where the complex transmission coefficients of each layer exhibit unit amplitude implying, thus, that each layer can be realized without utilizing resistive components (local design). We, instead, enlarge the set of feasible solutions of our design problem by allowing that the globally radiated power is preserved [18].
@@
The adopted channel model assumes frequency-flat block fading with path-loss-dependent variance. This simplified representation is intentionally used to isolate the effect of the artificial channel fluctuations introduced by the proposed ST-SIM architecture. In particular, the objective of this work is to evaluate how ST-SIM randomization enables opportunistic multiuser scheduling under limited CSIT, independently of propagation-induced randomness. More detailed propagation models including sparse multipath components, blockage effects, and spatial correlation across large apertures can be incorporated without modifying the proposed beamforming and scheduling framework. Likewise, recent studies have investigated SIMs in wideband communication systems, particularly MIMO systems based on orthogonal frequency-division multiplexing architectures where metasurface layers are optimized to jointly fit the frequency-selective channel response across subcarriers [33, 34]. The analysis presented here therefore focuses on the fundamental behavior of the ST-SIM mechanism rather than on environment-specific propagation characteristics.
@@ IV User scheduling with partial CSIT
and feeds the information $\text{SINR}_{u,n^{\star}_{u,m}}(m)$ back to the BS, for each time-slot. Based on the partial information provided by all the $U$ users in the network, the base station opportunistically schedules transmissions in every time-slot to the $N$ users who exhibit the highest $\text{SINR}_{u,n^{\star}_{u,m}}(m)$ values, where the maximization is performed over $u$ for each steering vector index $n\in\mathcal{N}$. Since each user feeds back only its preferred beam index $n_{u,m}^{\star}$ and the corresponding SINR value, each user can be associated with at most one beam in a given time-slot. Hence, distinct user assignment across beam indices is enforced by construction. The BS then groups the received feedback according to the reported beam index and, for each $n\in\mathcal{N}$, selects the user with the largest reported SINR among those declaring $n=n_{u,m}^{\star}$. If no user reports beam index $n$, that beam is left inactive during the corresponding slot. Therefore, fewer than $N$ users may be served in a slot, but no user is assigned to multiple beams simultaneously. In this manner, the system can opportunistically serve up to NMN\,M different users in each time interval $[0,T)$. The considered scheduling protocol assumes that, in each time slot, every user feeds back the index of its preferred steering vector together with a scalar quality metric (e.g., the corresponding SINR). Thus, the aggregate feedback overhead scales linearly with both the number of users and the number of time slots per coherence interval. Several lightweight reporting mechanisms can be adopted to further reduce the feedback load without modifying the proposed scheduling principle [38, 39]. For instance, users may report quantized versions of the SINR metric, employ threshold-based reporting (where feedback is transmitted only if the best-beam SINR exceeds a predefined threshold), or adopt contention-based or group-based feedback protocols suitable for dense-user scenarios.
We are now in the position of defining the user selected for each steering vector index. For a given time-slot $m\in\mathcal{M}$ and beam index $n\in\mathcal{N}$, let $\mathcal{U}_{n}(m)$ denote the set of users that reported $n=n_{u,m}^{\star}$. If $\mathcal{U}_{n}(m)\neq\emptyset$, the BS selects the user
$u_{n}^{\star}(m)\triangleq\arg\max_{u\in\mathcal{U}_{n}(m)}\text{SINR}_{u,n}(m).$ (34)
We denote by $\mathcal{U}^{\star}(m)$ the set of scheduled users in the $m$-th time-slot, i.e., $\mathcal{U}^{\star}(m)\subseteq\mathcal{U}$, with $|\mathcal{U}^{\star}(m)|\leq N$. If $\mathcal{U}_{n}(m)=\emptyset$, the corresponding beam $n$ remains inactive during that slot. For notational convenience, when a user is selected for beam $n$ within time-slot $m$, we denote the corresponding SINR as $\text{SINR}^{\star}_{u_{n}^{\star},n}(m)$. Given the received signal model (29), the achievable rate for user ${u}^{\star}_{n}$ in the $m$-th time-slot is
@@
where $\mathcal{R}_{{u}^{\star}_{n}}(m)$ denotes the achievable rate for user ${u}^{\star}_{n}$ in the $m$-th time-slot and is given by (35). The effective sum-rate accounts for the training and CSI-feedback overhead within each channel coherence interval. Specifically, we define
$\mathcal{R}_{\mathrm{eff}}\triangleq\xi\,\mathcal{R}$ (38)
where $0<\xi\leq 1$ denotes the fraction of symbols available for data transmission. The value of $\xi$ depends on the adopted transmission scheme, as well as on the associated training and CSI-feedback procedures.
@@
Remark 1: Unlike conventional S-only SIM schemes, the time modulation employed in the proposed ST-SIM architecture enables the exploitation of multiuser diversity. When the channel between the SIM and the users is slowly time-varying, with a coherence time $T$ spanning multiple symbol periods (i.e., $T=M\,T_{s}=M\,P\,T_{\text{b}}$, with $M\,P\gg 1$), the channel remains nearly constant over several transmission intervals. In such a scenario, only $N$ out of the $U$ users can be scheduled and effectively served within each coherence interval. As the total number of users $U$ increases, this leads to significant unfairness across the network [40]. In contrast, the proposed ST-SIM-based beamforming approach introduces artificial time variations into the wireless channel via time modulation. This temporal diversity allows the system to opportunistically schedule up to $MN$ users for transmission. Consequently, the ST-SIM architecture improves fairness and enables the exploitation of multiuser diversity, even under slow channel dynamics.
@@ V Transmission Coefficient Synthesis for randomized ST-SIM
At the beginning of each time slot, the ST-SIM has to generate a set of $N$ beamforming vectors $\{\widetilde{\boldsymbol{g}}_{n}(t)\}_{n=0}^{N-1}$, each used to modulate the corresponding information signal $d_{n}(t)$. Recalling the structure of the overall forward propagation matrix in (15), the generation of these steering vectors involves two distinct steps. The first step consists of generating a fixed matrix ${\mathbf{G}}_{0}$, which models the forward propagation through the S-only multilayer block of the SIM. This step is performed once, at the beginning of each channel coherence interval of duration $T$. The second step, in contrast, is repeated at the beginning of each time-slot of duration $T_{\text{s}}$ and consists of randomly generating the rapidly TV transmission coefficients $\boldsymbol{\delta}(t)$ of the initial ST DAL.
Step 1 - Generation of the matrix ${\mathbf{G}}_{0}$: Let ${\mathbf{G}}_{\text{targ}}\in\mathbb{C}^{V\times Z}$ denote the target matrix that models the forward propagation through the S-only multilayer block of the SIM, namely, from the output of the ST DAL to the output of the S-only DAL. It can be expressed as
@@
where ${\mathbf{g}}_{\text{targ},n}\in\mathbb{C}^{V}$ denotes the desired steering vector associated with the $n$-th transmitted stream. The optimization framework does not require any particular structure of ${\mathbf{G}}_{\text{targ}}$, whose columns only need to satisfy the norm constraint (25).
Our objective is to synthesize the transmission coefficients $\{\boldsymbol{\gamma}_{\ell}\}_{\ell\in\mathcal{L}^{(\text{s})}}$ of the S-only metasurface layers such that the matrix ${\mathbf{G}}_{0}$ in (15) closely approximates ${\mathbf{G}}_{\text{targ}}$. This leads to the constrained least-squares (LS) optimization problem
@@
$\|{\mathbf{g}}_{z}^{(0)}\|^{2}=\frac{N}{\beta^{2}\,\|{\mathbf{W}}_{1}\|^{2}},\quad z\in\mathcal{Z}$ (46)
@@
The PGD iterations (51)-(52) continue until a convergence criterion is met, either a sufficiently small variation in the cost function or a maximum number of iterations $\kappa_{\max}$ is reached. Upon convergence, the set of optimized transmission coefficients $\{\boldsymbol{\gamma}^{\star}_{\ell}\}_{\ell\in\mathcal{L}^{(\text{s})}}$ configures the S-only part of the SIM.
@@
with ${\mathbf{1}}_{Q}$ denoting the $Q\times 1$ all-ones vector, and $\circ$ being the Hadamard (element-wise) product. The per-iteration complexity of the PGD synthesis is dominated by the computation of the matrices $\mathbf{E}_{\ell}^{H}\mathbf{E}_{\ell}$ and $\mathbf{B}_{\ell}^{\ast}\mathbf{B}_{\ell}^{T}$, yielding a complexity scaling of $\mathcal{O}\left((L-1)\,(V+Z)\,Q^{2}\right)$ per iteration.